# Universality theorems for zeros of random real polynomials with fixed coefficients

### v1

@@ Abstract
Consider a monic polynomial of degree $n$ whose subleading coefficients are independent, identically distributed, nondegenerate random variables having zero mean, unit variance, and finite moments of all orders, and let $m\geq 0$ be a fixed integer. We prove that such a random monic polynomial has exactly $m$ real zeros with probability $n^{-b+o(1)}$ as $n\to\infty$ through integers of the same parity as $m$, where $b\approx 0.76$ is a positive constant. More generally, we determine conditions under which a similar asymptotic formula describes the corresponding probability for families of random real polynomials with multiple fixed coefficients. Our work extends well-known universality results of Dembo, Poonen, Shao, and Zeitouni, whose considered the family of real polynomials with all coefficients random.
As a number-theoretic consequence of these results, we deduce that an algebraic integer $\alpha$ of degree $n$ has exactly $m$ real Galois conjugates with probability $n^{-b+o(1)}$, when such $\alpha$ are ordered by the heights of their minimal polynomials.
@@ 1 Introduction
In [DPSZ02], Dembo, Poonen, Shao, and Zeitouni (henceforth, DPSZ) study the asymptotic behavior of $P_{n}^{*}$ in the limit as $n\to\infty$. Strikingly, they prove that there exists a universal positive constant $b$, independent of $n$, such that $P_{n}^{*}=n^{-b+o(1)}$ (i.e., the limit $\lim_{n\to\infty}\log_{n}P_{n}^{*}$ exists and equals $-b$); see Theorem 1.1(a) in loc. cit. While the value of $b$ remains unknown, it was proven by DPSZ that $0.4\leq b\leq 2$, and their numerical simulations suggest that $b\approx 0.76$. In light of their remarkable result, it is natural to ask the following questions:
@@
    Does the above result of DPSZ for the full family of random polynomials $f_{n}^{*}(x)$ admit an analogue for the special family of monic random polynomials $f_{n}(x)$? More specifically, is there a constant $b^{\prime}>0$ such that $P_{n}=n^{-b^{\prime}+o(1)}$?
@@
    If the answer to the first question is yes, then how is $b^{\prime}$ related to $b$? Are they equal?
@@
In this setting, one can ask how $P_{n,j}$ and $P_{n,j}^{*}$ behave asymptotically as $n\to\infty$. When $j=0$, we have that $P_{n,0}=P_{n}$ and $P_{n,0}^{*}=P_{n}^{*}$, and the aforementioned asymptotic formulas apply. In [DPSZ02], DPSZ prove that, for the full family of random polynomials $f_{n}(x)$, the same asymptotic formula holds — i.e., we have that $P_{n,j}^{*}=n^{-b+o(1)}$ — for each $j$, all the way up to $j=o(\log n/\log\log n)$; see Theorem 1.2 in loc. cit. Our next result gives the analogue for the family of random monic polynomials $f_{n}^{*}(x)$:
@@
Let $j\in\mathbb{N}$ be such that $j\equiv n-1\pmod{2}$. Then we have that $P_{n,j}=n^{-b+o(1)}$ as $n\to\infty$. In fact, $f_{n}(x)$ has at most $o(\log n/\log\log n)$ real zeros with probability $n^{-b+o(1)}$.
@@ 1.1 Generalizations of Theorems 1 and 2 to polynomials with multiple fixed coefficients
One cannot reasonably expect that an unconditional analogue of Theorem 1 would hold with $P_{n}$ replaced by, say, $P_{n,S,0}$ (where by the subscript “00” we mean that $\gamma_{n}\equiv 0$ is taken to be identically zero for each $n$). Indeed, one can choose the coefficient data — i.e., the data of the set $S$, the fixed coefficients $c_{i}$ for $i\in S$, and the distribution of the random coefficients $a_{i}$ — in such a way that $P_{n,S,0}=0$ for infinitely many $n$ [^1] As we detail in Example 15 (to follow), one way to do this is to choose the distribution of the $a_{i}$ to be bounded in absolute value by some constant $C$, and to choose some of the $c_{i}$ to be considerably less than $-C$. We then analyze the positivity of $f_{n,S}(x)$ on three ranges of $|x|$. First, when $|x|<1$, the fixed terms matter little, and the lower-degree random terms dominate, making $f_{n,S}(x)>0$ with positive probability. The same occurs when $|x|-1$ is sufficiently large, in which case the leading term is positive and dominates over all other terms. But in the middle range, when $10\leq|x|-$ 1 is sufficiently small, the fixed terms can contribute large negative quantities to the value of $f_{n,S}(x)$, forcing it to be negative regardless of the values of the random coefficients.
@@
We have the upper bound $P_{n,S,\gamma_{n}}\leq n^{-b+o(1)}$. Furthermore, if the chosen data are nice, then we have equality $P_{n,S,\gamma_{n}}=n^{-b+o(1)}$.
@@
as $n\to\infty$. Then we have that $P_{n,S,\gamma_{n}}=n^{-b+o(1)}$.
@@
Let $j\in\mathbb{N}$ be such that $j\equiv n-1\pmod{2}$. Then we have the upper bound $P_{n,j,S}\leq n^{-b+o(1)}$. If the chosen data are nice, then we have equality $P_{n,j,S}=n^{-b+o(1)}$. Furthermore, the polynomial $f_{n}(x)$ has at most $o(\log n/\log\log n)$ real zeros with probability at most $n^{-b+o(1)}$, with equality if the coefficient data are nice.
@@
Let notation be as above, and suppose that $k(n)=n^{o(1)}$ in the case of Gaussian coefficients, and that $k(n)=n^{o(1/\sqrt{\log n})}$ in the case of general coefficients. Then we have $P_{n,S_{n},\gamma_{n}}=n^{-b+o(1)}$.
@@
Let notation be as above, and suppose that $k(n)=o(\log n)$. Let $j\in\mathbb{N}$ be such that $j\equiv n-1\pmod{2}$. Then we have $P_{n,j,S_{n}}=n^{-b+o(1)}$.
@@ 1.2 Application to counting algebraic integers
Let $j\in\mathbb{N}$ be fixed. When algebraic integers $\alpha$ of degree $n\equiv j\pmod{2}$ are ordered by height, the density of $\alpha$ that are $j$-realizable is $n^{-b+o(1)}$.
@@
The problem of counting algebraic integers satisfying interesting conditions at the archimedean places has been studied before in the literature. Of particular relevance to the present paper is the work of Calegari and Huang [CH17] (cf. the closely related earlier work of Akiyama and Pethő [AP14a, AP14b]), who consider algebraic integers ordered by a somewhat different height function, namely, the absolute value of the largest Galois conjugate. With respect to this height, they determine precise asymptotics for densities of various types of algebraic integers, including the totally complex subfamily, for which they obtain a density $\asymp n^{-3/8}$. Proving this result amounts to studying monic real polynomials, all of whose roots have absolute value at most $1$; note that such polynomials are special in that they are positive on $\mathbb{R}$ if and only if they are positive on $[-1,1]$. On the other hand, by modifying the proof of Theorem 1, it is possible to show that a random monic polynomial is positive on $[-1,1]$ with probability $n^{-b/2+o(1)}$. As observed in [CH17, Remark 1.5], this raises the intriguing question of whether the densities $n^{-3/8}$ and $n^{-b/2+o(1)}$ agree — i.e., whether $b$ is in fact equal to $3/4$, a result that would seemingly concur with the aforementioned numerical simulations performed by DPSZ, which indicated that $b\approx 0.76$.
@@ 4.1.2 The case where $n-1,j$ are odd
Since the $a_{i}$ are of zero mean and unit variance, there exist $\beta<0<\alpha$ such that $\alpha,\beta\in\mathcal{V}$. We may assume without loss of generality that $\alpha>-\beta$. Let $s\geq 4$ be an even integer such that $\alpha+(s-1)\beta<0$, and define polynomials
@@ Acknowledgments
We thank Kumar Murty, Bjorn Poonen, Qi-Man Shao, and Melanie Matchett Wood for helpful conversations. AS was supported by the National Science Foundation, under the Graduate Research Fellowship, as well as Award No. 2202839.

### v2

@@ Abstract
Consider a monic polynomial of degree $n$ whose subleading coefficients are independent, identically distributed, nondegenerate random variables having zero mean, unit variance, and finite moments of all orders, and let $m\geq 0$ be a fixed integer. We prove that such a random monic polynomial has exactly $m$ real zeros with probability $n^{-3/4+o(1)}$ as $n\to\infty$ through integers of the same parity as $m$. More generally, we determine conditions under which a similar asymptotic formula describes the corresponding probability for families of random real polynomials with multiple fixed coefficients. Our work extends well-known universality results of Dembo, Poonen, Shao, and Zeitouni, who considered the family of real polynomials with all coefficients random.
As a number-theoretic consequence of these results, we deduce that an algebraic integer $\alpha$ of degree $n$ has exactly $m$ real Galois conjugates with probability $n^{-3/4+o(1)}$, when such $\alpha$ are ordered by the heights of their minimal polynomials.
@@ 1 Introduction
In [DPSZ02], Dembo, Poonen, Shao, and Zeitouni (henceforth, DPSZ) studied the asymptotic behavior of $P_{n}^{*}$ in the limit as $n\to\infty$. Strikingly, they proved that there exists a universal positive constant $b$, independent of $n$, such that $P_{n}^{*}=n^{-b+o(1)}$ (i.e., the limit $\lim_{n\to\infty}\log_{n}P_{n}^{*}$ exists and equals $-b$); see Theorem 1.1(a) in loc. cit. They further showed that $0.4\leq b\leq 2$ and performed numerical simulations indicating that $b\approx 0.76$. It was later determined in works of Poplavskyi and Schehr [PS18, Equation (4)] and FitzGerald, Tribe, and Zaboronski [FTZ22, Proposition 5] that $b=3/4$; for further discussion of the literature on the constant $b$, see Remark 13 (to follow).
In light of these remarkable results, it is natural to ask the following pair of questions:
@@
    Does the above asymptotic formula for the full family of random polynomials $f_{n}^{*}(x)$ admit an analogue for the special family of monic random polynomials $f_{n}(x)$? More specifically, is there a constant $b^{\prime}>0$ such that $P_{n}=n^{-b^{\prime}+o(1)}$?
@@
    If the answer to the first question is yes, then how is $b^{\prime}$ related to $b=3/4$? Are they equal?
@@
In this setting, one can ask how $P_{n,j}$ and $P_{n,j}^{*}$ behave asymptotically as $n\to\infty$. When $j=0$, we have that $P_{n,0}=P_{n}$ and $P_{n,0}^{*}=P_{n}^{*}$, and the aforementioned asymptotic formulas apply. In [DPSZ02], DPSZ prove that, for the full family of random polynomials $f_{n}(x)$, the same asymptotic formula holds — i.e., we have that $P_{n,j}^{*}=n^{-3/4+o(1)}$ — for each $j$, all the way up to $j=o(\log n/\log\log n)$; see Theorem 1.2 in loc. cit. Our next result gives the analogue for the family of random monic polynomials $f_{n}^{*}(x)$:
@@
Let $j\in\mathbb{N}$ be such that $j\equiv n-1\pmod{2}$. Then we have that $P_{n,j}=n^{-3/4+o(1)}$ as $n\to\infty$. In fact, $f_{n}(x)$ has at most $o(\log n/\log\log n)$ real zeros with probability $n^{-3/4+o(1)}$.
@@ 1.1 Generalizations of Theorems 1 and 2 to polynomials with multiple fixed coefficients
One cannot reasonably expect that an unconditional analogue of Theorem 1 would hold with $P_{n}$ replaced by, say, $P_{n,S,0}$ (where by the subscript “00” we mean that $\gamma_{n}\equiv 0$ is taken to be identically zero for each $n$). Indeed, one can choose the coefficient data — i.e., the data of the set $S$, the fixed coefficients $c_{i}$ for $i\in S$, and the distribution of the random coefficients $a_{i}$ — in such a way that $P_{n,S,0}=0$ for infinitely many $n$. As we detail in Example 16 (to follow), one way to do this is to choose the distribution of the $a_{i}$ to be bounded in absolute value by some constant $C$, and to choose some of the $c_{i}$ to be considerably less than $-C$. We then analyze the positivity of $f_{n,S}(x)$ on three ranges of $|x|$. First, when $|x|<1$, the fixed terms matter little, and the lower-degree random terms dominate, making $f_{n,S}(x)>0$ with positive probability. The same occurs when $|x|-1$ is sufficiently large, in which case the leading term is positive and dominates over all other terms. But in the middle range, when $10\leq|x|-$ 1 is sufficiently small, the fixed terms can contribute large negative quantities to the value of $f_{n,S}(x)$, forcing it to be negative regardless of the values of the random coefficients.
@@
We have the upper bound $P_{n,S,\gamma_{n}}\leq n^{-3/4+o(1)}$. Furthermore, if the chosen data are nice, then we have equality $P_{n,S,\gamma_{n}}=n^{-3/4+o(1)}$.
@@
as $n\to\infty$. Then we have that $P_{n,S,\gamma_{n}}=n^{-3/4+o(1)}$.
@@
Let $j\in\mathbb{N}$ be such that $j\equiv n-1\pmod{2}$. Then we have the upper bound $P_{n,j,S}\leq n^{-3/4+o(1)}$. If the chosen data are nice, then we have equality $P_{n,j,S}=n^{-3/4+o(1)}$. Furthermore, the polynomial $f_{n}(x)$ has at most $o(\log n/\log\log n)$ real zeros with probability at most $n^{-3/4+o(1)}$, with equality if the coefficient data are nice.
@@
Let notation be as above, and suppose that $k(n)=n^{o(1)}$ in the case of Gaussian coefficients, and that $k(n)=n^{o(1/\sqrt{\log n})}$ in the case of general coefficients. Then we have $P_{n,S_{n},\gamma_{n}}=n^{-3/4+o(1)}$.
@@
Let notation be as above, and suppose that $k(n)=o(\log n)$. Let $j\in\mathbb{N}$ be such that $j\equiv n-1\pmod{2}$. Then we have $P_{n,j,S_{n}}=n^{-3/4+o(1)}$.
@@ 1.2 Application to counting algebraic integers
Let $j\in\mathbb{N}$ be fixed. When algebraic integers $\alpha$ of degree $n\equiv j\pmod{2}$ are ordered by height, the density of $\alpha$ that are $j$-realizable is $n^{-3/4+o(1)}$.
@@
The problem of counting algebraic integers satisfying interesting conditions at the archimedean places has been studied before in the literature. Of particular relevance to the present paper is the work of Calegari and Huang [CH17] (cf. the closely related earlier work of Akiyama and Pethő [AP14a, AP14b]), who consider algebraic integers ordered by a somewhat different height function, namely, the absolute value of the largest Galois conjugate. With respect to this height, they determine precise asymptotics for densities of various types of algebraic integers, including the totally complex subfamily, for which they obtain a density $\asymp n^{-3/8}$. Proving this result amounts to studying monic real polynomials, all of whose roots have absolute value at most $1$; note that such polynomials are special in that they are positive on $\mathbb{R}$ if and only if they are positive on $[-1,1]$. We note that, by modifying the proof of Theorem 1, it is possible to show that a random monic polynomial is positive on $[-1,1]$ with probability $n^{-3/8+o(1)}$.
@@ 1.3 Summary of related earlier work
Remark 13.
The problem of determining both the physical meaning and the precise value of the constant $b$ introduced by DPSZ has garnered significant attention in the literature. Years after DPSZ, Schehr and Majumdar [SM07, SM08] proved that $b$ is the so-called persistence exponent of the two-dimensional heat equation. Other works, including those of Li and Shao [LS02] and Molchan [Mol12] improved on the bounds on $b$ obtained by DPSZ. Meanwhile, in the same direction as some of the work of Schehr and Majumdar, Dembo and Mukherjee generalized the results of DPSZ to the case of independent coefficients with polynomially growing variance [DM15]. Despite studying a different family of random polynomials, Calegari and Huang were motivated, after their aforementioned result, to ask whether $b=3/4$ [CH17]. Finally, almost two decades after DPSZ, Poplavskyi and Schehr [PS18, Equation (4)] used the truncated real ensemble of orthogonal matrices to show that $b=3/4$. More recently, FitzGerald, Tribe, and Zaboronski [FTZ22, Proposition 5] gave a different derivation of the value of $b$, using certain asymptotic results for Fredholm Pfaffians. For the sake of clarity, we choose to write $b$ for the constant in the rest of the paper, even though its value is known to be $3/4$.
@@ 4.1.2 The case where $n-1,j$ are odd
Since the $a_{i}$ are of zero mean and unit variance, there exist $\alpha<0<\beta$ such that $\alpha,\beta\in\mathcal{V}$. We may assume without loss of generality that $-\alpha>\beta$. Let $s\geq 4$ be an even integer such that $\alpha+(s-1)\beta<0$, and define polynomials
@@ Acknowledgments
We are grateful to Kumar Murty, Bjorn Poonen, Qi-Man Shao, and Melanie Matchett Wood for helpful conversations. We also thank Will FitzGerald for pointing us to the references [FTZ22] and [PS18]. AS was supported by the National Science Foundation, under the Graduate Research Fellowship, as well as Award No. 2202839.